OpenStudy (anonymous):

cos(2x-pi/4)=sqrt2/2

OpenStudy (anonymous):

substitute 2x-pi/4 for "u" and solve it like that first and then plug it back in to that answer

OpenStudy (anonymous):

\[\Large \cos(2x-\frac{\pi}{4})=\frac{\sqrt{2}}{2}\] correct?

OpenStudy (anonymous):

\[ \Large 2x-\frac{\pi}{4}=\pm \frac \pi 4 + 2 k \pi\\ \]

OpenStudy (anonymous):

You can finish it now.